# Peeling Context from Cause for Multimodal Molecular Property Prediction

### v1

@@ Appendix B Additional Experiments
B.1 Additional Ablation Study on Framework Design
Table 4: Additional ablation study on framework design across four datasets.
Setting MAE MSE $R^{2}$
ESOL
CLaP $0.4456\pm 0.0151$ $0.3583\pm 0.0095$ $0.9157\pm 0.0022$
Average causal layer $0.4798\pm 0.0172$ $0.4377\pm 0.0064$ $0.8990\pm 0.0218$
No mono penalty $0.5692\pm 0.0192$ $0.5740\pm 0.0606$ $0.8600\pm 0.0165$
FreeSolv
CLaP $0.7020\pm 0.0044$ $0.8866\pm 0.0522$ $0.9160\pm 0.0049$
Average causal layer $0.9612\pm 0.0239$ $1.5843\pm 0.0435$ $0.8498\pm 0.0420$
No mono penalty $1.0241\pm 0.1422$ $1.7210\pm 0.3891$ $0.8368\pm 0.0369$
Lipo
CLaP $0.4672\pm 0.0020$ $0.3645\pm 0.0274$ $0.7515\pm 0.0187$
Average causal layer $0.5621\pm 0.0124$ $0.5041\pm 0.0156$ $0.6563\pm 0.0016$
No mono penalty $0.5288\pm 0.0591$ $0.4590\pm 0.0845$ $0.6870\pm 0.0576$
CycPeptMPDB
CLaP $0.3056\pm 0.0080$ $0.1644\pm 0.0082$ $0.7282\pm 0.0136$
Average causal layer $0.3227\pm 0.0118$ $0.1859\pm 0.0067$ $0.6943\pm 0.0160$
No mono penalty $0.3137\pm 0.0051$ $0.1803\pm 0.0032$ $0.7051\pm 0.0089$
@@ B.1 Additional Ablation Study on Framework Design
Average causal across layers. This ablation probes the depthwise correlation design. The targets $\{\rho_{\ell}\}$ increase with depth (Eq. equation 7) and, together with $\mathcal{L}_{\mathrm{corr}}$ and $\mathcal{L}_{\mathrm{mono}}$ (Eqs. equation 8, equation 9), train the deepest causal block to become progressively more label–relevant. Here we replace the final-layer causal output with a uniform average of $\{{\bm{c}}^{(\ell)}\}_{\ell=1}^{L}$. Given $\mathrm{corr}_{1}\!\leq\!\cdots\!\leq\!\mathrm{corr}_{L}$ under $\mathcal{L}_{\mathrm{corr}}+\mathcal{L}_{\mathrm{mono}}$, averaging $\bar{c}=\tfrac{1}{L}\sum_{\ell}{\bm{c}}^{(\ell)}$ mixes in low-correlation early representations. For centered variables, $\mathrm{Corr}(\bar{c},y)\leq\mathrm{Corr}({\bm{c}}^{(L)},y)$, with strict inequality unless all ${\bm{c}}^{(\ell)}$ are colinear with ${\bm{c}}^{(L)}$. Thus using ${\bm{c}}^{(L)}$ is theoretically favored.
@@
In our experiments, averaging across layers undermines depthwise specialization and mixes early-layer representations that retain substantial context information into the final predictor, contaminating the causal signal and degrading performance (e.g., $R^{2}:0.728\!\to\!0.694$, MSE: $0.164\!\to\!0.185$). These results validate the design: the last causal layer concentrates the causal signal, while earlier layers serve as progressively coarser precursors.
@@
Together, the correlation schedule shapes what each depth should capture; the monotonicity term enforces how the hierarchy evolves; the trivial branch determines where non-causal residuals go.
B.2 Multi-modality Design
Table 5: Effect of causal peeling under unimodal vs. multimodal settings across four datasets.
Setting MAE MSE $R^{2}$
ESOL
Multimodal
   w causal-trivial split 0.4456 $\pm$ 0.0151 0.3583 $\pm$ 0.0095 0.9157 $\pm$ 0.0022
   w/o causal-trivial split 0.5094 $\pm$ 0.0528 0.4939 $\pm$ 0.1034 0.8838 $\pm$ 0.0243
2D Graph
   w causal-trivial split 0.9453 $\pm$ 0.0156 1.5993 $\pm$ 0.0719 0.6238 $\pm$ 0.0169
   w/o causal-trivial split 0.9631 $\pm$ 0.0092 1.6931 $\pm$ 0.0817 0.6018 $\pm$ 0.0192
3D Geometry
   w causal-trivial split 0.5322 $\pm$ 0.0029 0.5235 $\pm$ 0.0268 0.8769 $\pm$ 0.0063
   w/o causal-trivial split 0.5219 $\pm$ 0.0055 0.5036 $\pm$ 0.0088 0.8816 $\pm$ 0.0021
FreeSolv
Multimodal
   w causal-trivial split 0.7020 $\pm$ 0.0044 0.8866 $\pm$ 0.0522 0.9160 $\pm$ 0.0049
   w/o causal-trivial split 1.2858 $\pm$ 0.2120 2.5719 $\pm$ 0.8194 0.7562 $\pm$ 0.0777
2D Graph
   w causal-trivial split 1.1267 $\pm$ 0.0098 1.9876 $\pm$ 0.2791 0.8116 $\pm$ 0.0265
   w/o causal-trivial split 1.2003 $\pm$ 0.0410 2.4157 $\pm$ 0.2852 0.7710 $\pm$ 0.0270
3D Geometry
   w causal-trivial split 1.0640 $\pm$ 0.0001 1.8055 $\pm$ 0.0591 0.8288 $\pm$ 0.0056
   w/o causal-trivial split 1.0721 $\pm$ 0.0875 1.9262 $\pm$ 0.2170 0.8174 $\pm$ 0.0206
Lipo
Multimodal
   w causal-trivial split 0.4672 $\pm$ 0.0021 0.3645 $\pm$ 0.0274 0.7515 $\pm$ 0.0187
   w/o causal-trivial split 0.5145 $\pm$ 0.0425 0.4302 $\pm$ 0.0442 0.7067 $\pm$ 0.0381
2D Graph
   w causal-trivial split 0.5171 $\pm$ 0.0234 0.4506 $\pm$ 0.0424 0.6928 $\pm$ 0.0289
   w/o causal-trivial split 0.5230 $\pm$ 0.0283 0.4504 $\pm$ 0.0411 0.6929 $\pm$ 0.0280
3D Geometry
   w causal-trivial split 0.6412 $\pm$ 0.0339 0.6519 $\pm$ 0.0784 0.5556 $\pm$ 0.0534
   w/o causal-trivial split 0.6260 $\pm$ 0.0739 0.6449 $\pm$ 0.1567 0.5603 $\pm$ 0.1069
CycPeptMPDB
Multimodal
   w causal-trivial split 0.3056 $\pm$ 0.0080 0.1644 $\pm$ 0.0082 0.7282 $\pm$ 0.0136
   w/o causal-trivial split 0.3120 $\pm$ 0.0047 0.1834 $\pm$ 0.0063 0.6969 $\pm$ 0.0104
2D Graph
   w causal-trivial split 0.4068 $\pm$ 0.0001 0.2889 $\pm$ 0.0112 0.5226 $\pm$ 0.0184
   w/o causal-trivial split 0.4077 $\pm$ 0.0069 0.2884 $\pm$ 0.0069 0.5234 $\pm$ 0.0114
HELM Notation
   w causal-trivial split 0.3310 $\pm$ 0.0034 0.2023 $\pm$ 0.0005 0.6657 $\pm$ 0.0008
   w/o causal-trivial split 0.3194 $\pm$ 0.0026 0.1910 $\pm$ 0.0025 0.6843 $\pm$ 0.0042
3D Geometry
   w causal-trivial split 0.4221 $\pm$ 0.0111 0.3102 $\pm$ 0.0164 0.4874 $\pm$ 0.0271
   w/o causal-trivial split 0.4244 $\pm$ 0.0087 0.3071 $\pm$ 0.0134 0.4924 $\pm$ 0.0222
In this section, we assess how the causal–trivial split behaves under different modality configurations. CycPeptMPDB provides three complementary views per sample (SM, PE, GE), whereas ESOL, FreeSolv, and Lipo provide two views (SM and GE). Table 5 compares unimodal variants with multimodal fusion, each trained with and without the split.
With two or more modalities, enabling the split generally yields better fit and stability across datasets. For instance, MSE decreases on CycPeptMPDB from 0.183 to 0.164 and on ESOL from 0.494 to 0.358, with similar gains on FreeSolv and Lipo. This pattern aligns with Sec. 2.2. After node and edge splitting, the depthwise gate ${\bm{w}}^{(\ell)}$ emphasizes the modality whose causal branch carries batch invariant evidence. The correlation schedule aligns each layer’s causal scalar with the label, and the residual path learns remaining context with a stop gradient to prevent leakage into the causal branch. Fusion supplies complementary cues that the curriculum can promote and confounding that it can peel away, which strengthens accuracy.
For single view variants, changes are small or mixed. ESOL based on 2D shows a modest $R^{2}$ gain from 0.602 to 0.624, while ESOL based on 3D declines slightly from 0.882 to 0.877. Lipo based on 2D is essentially unchanged at about 0.693, whereas Lipo based on 3D decreases a bit from 0.560 to 0.556. On CycPeptMPDB, 2D and 3D are near ties and HELM drops from 0.684 to 0.666. FreeSolv is an exception where single view 2D and 3D both improve to 0.812 and 0.829. In practice, single view deployments use a milder target schedule or omit the split while retaining residual calibration. The split functions best as a fusion amplifier when depthwise correlation guidance and cross modal gating can reinforce each other.
Input : multi-view data $(x_{\mathrm{SM}},x_{\mathrm{PE}},x_{\mathrm{GE}})$ with label $y$; depth $L$; epochs $E$; batch size $B$; correlation targets $\{\rho_{\ell}\}$; weights $\lambda_{\mathrm{caus}},\lambda_{\mathrm{mono}},\lambda_{\mathrm{unif}}$.
Output : trained predictor $\widehat{y}=c^{\star}+t$.
for $e=1$ to $E$ do
 for each mini-batch do
      Encode inputs into embeddings.
    for $\ell=1$ to $L$ do
         Split into causal/trivial parts, fuse across modalities, get scalars ${\bm{c}}^{(\ell)},{\bm{t}}^{(\ell)}$.
    $c^{\star}\leftarrow{\bm{c}}^{(L)}$,  $t\leftarrow\sum_{\ell}{\bm{t}}^{(L)}$,  $\widehat{y}\leftarrow c^{\star}+t$.
      Compute losses: $\mathcal{L}_{\mathrm{pred}}$ (MSE), $\mathcal{L}_{\mathrm{corr}}$ (correlation), $\mathcal{L}_{\mathrm{mono}}$ (monotonicity), $\mathcal{L}_{\mathrm{triv}}$ (residual).
      Total: $\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{pred}}+\lambda_{\mathrm{caus}}\mathcal{L}_{\mathrm{corr}}+\lambda_{\mathrm{mono}}\mathcal{L}_{\mathrm{mono}}+\lambda_{\mathrm{unif}}\mathcal{L}_{\mathrm{triv}}$;
      optimizer step w.r.t. $\nabla\mathcal{L}_{\mathrm{total}}$.
Algorithm 1 Layer-wise causal–trivial peeling
B.3 Cross-Modal Consistency
Table 6: Effect of consistency weight across datasets.
Setting MAE MSE $R^{2}$
ESOL
CLaP $0.4456\pm 0.0151$ $0.3583\pm 0.0095$ $0.9157\pm 0.0022$
$\lambda_{\text{cons}}$ =0.5=0.5 $0.5242\pm 0.0618$ $0.5106\pm 0.1179$ $0.8799\pm 0.0277$
$\lambda_{\text{cons}}$ =1.0=1.0 $192\pm 0.0439 0.505$$050\pm 0.0583$ $0.8812\pm 0.0137$
FreeSolv
CLaP $0.7020\pm 0.0044$ $0.8866\pm 0.0522$ $0.9160\pm 0.0049$
$\lambda_{\text{cons}}$ =0.5=0.5 $1.0550\pm 0.3457$ $1.8375\pm 1.0257$ $0.8258\pm 0.0972$
$\lambda_{\text{cons}}$ =1.0=1.0 $1.1412\pm 0.2175$ $2.0513\pm 0.7842$ $0.8055\pm 0.0743$
Lipo
CLaP $0.4672\pm 0.0201$ $0.3645\pm 0.0274$ $0.7515\pm 0.0187$
$\lambda_{\text{cons}}$ =0.5=0.5 $0.5649\pm 0.0272$ $0.5002\pm 0.0496$ $0.6899\pm 0.0338$
$\lambda_{\text{cons}}$ =1.0=1.0 $0.5597\pm 0.0920$ $0.4949\pm 0.1213$ $0.6626\pm 0.0827$
CycPeptMPDB
CLaP $0.3056\pm 0.0080$ $0.1644\pm 0.0082$ $0.7282\pm 0.0136$
$\lambda_{\text{cons}}$ =0.5=0.5 $0.3204\pm 0.0002$ $0.1934\pm 0.0097$ $0.6803\pm 0.0160$
$\lambda_{\text{cons}}$ =1.0=1.0 $0.3154\pm 0.0015$ $0.1839\pm 0.0012$ $0.6960\pm 0.0020$
We vary a cosine-similarity penalty between modality-specific embeddings with weight $\lambda_{\text{cons}}$ to test whether enforcing cross-modal agreement helps or hurts (Table 6). Across datasets, adding this penalty degrades the metrics compared to our default, with larger weights producing larger drops. The effect is most visible on FreeSolv and Lipo, and still present on CycPeptMPDB and ESOL.
This pattern is consistent with our design in Sec. 2.2. The model benefits from complementarity across modalities: the split routes label-relevant signal to the causal branch while the fusion gate selects the modality that carries the most batch-invariant evidence at each depth. Forcing embeddings to be too similar collapses modality-specific subspaces, reduces the gate’s effective choice set, and makes batch-coupled shortcuts more likely to be shared. The correlation curriculum then has less diverse causal evidence to promote, and the residual path absorbs less of the confounding context. In practice, we keep $\lambda_{\text{cons}}$ at zero or very small to preserve complementarity.
B.4 Peeling Depth on Additional Datasets
Table 7: Peeling depth ablation on ESOL, FreeSolv, and Lipo.
Layer MAE MSE $R^{2}$
ESOL
1 $0.4879\pm 0.0098$ $0.4293\pm 0.0101$ $0.8990\pm 0.0024$
2 $\bm{0.4456\pm 0.0151}$ $\bm{0.3583\pm 0.0095}$ $\bm{0.9157\pm 0.0022}$
3 $0.5495\pm 0.1119$ $0.5794\pm 0.2043$ $0.8637\pm 0.0565$
4 $0.5463\pm 0.0235$ $0.5335\pm 0.0234$ $0.8745\pm 0.0055$
FreeSolv
1 $0.9733\pm 0.1168$ $1.9424\pm 0.1979$ $0.8159\pm 0.0188$
2 $\bm{0.7020\pm 0.0044}$ $\bm{0.8866\pm 0.0522}$ $\bm{0.9160\pm 0.0049}$
3 $0.8987\pm 0.0910$ $1.4897\pm 0.3848$ $0.8588\pm 0.0365$
4 $0.9849\pm 0.0834$ $1.5328\pm 1.1299$ $0.8547\pm 0.0123$
Lipo
1 $0.5201\pm 0.0347$ $0.4506\pm 0.0626$ $0.6928\pm 0.0427$
2 $0.5079\pm 0.0504$ $0.4356\pm 0.0676$ $0.7030\pm 0.0461$
3 $\bm{0.4672\pm 0.0201}$ $\bm{0.3645\pm 0.0274}$ $\bm{0.7515\pm 0.0187}$
4 $0.5376\pm 0.0366$ $0.4699\pm 0.0455$ $0.6796\pm 0.0310$
We extend the depth sweep to the remaining datasets. Table 7 shows a consistent U-shaped trend as the number of causal blocks increases. With too few layers the hierarchy under-peels. As depth grows beyond a dataset-dependent sweet spot the hierarchy over-peels.
ESOL and FreeSolv favor shallow hierarchies (two causal blocks), whereas Lipo benefits from one additional block, likely due to greater dataset complexity. This pattern is consistent with bias–variance tradeoffs and batch-wise invariance objective. The additional layer is useful when it tries to capture more complex signal.
In practice, we select depth by monitoring the layerwise correlation curve against the target schedule and choosing the smallest $L$ at which the validation error reaches its first stable minimum while correlations remain monotone and close to the anchors. Beyond that point additional layers yield diminishing returns and increase sensitivity to context.
@@ C.1 ESOL Dataset
Figure 6: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s water affinity the most.
At the optimal depth (L2), CLaP assigns the highest weights to chloro substituents and adjacent aryl carbons that cluster via short Cl–C/Cl–Cl distances in 3D; the fused readout preserves this through-space signal while the 2D branch highlights halogenated positions from connectivity alone. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (a single hydrophobic patch) is retained. The outcome matches physical chemistry: contiguous chlorination increases non-polar surface and reduces hydrogen-bonding capacity, driving logS lower. Under-peeling spreads attribution across the ring (shortcuts), and over-peeling reintroduces residual context, illustrating why CLaP’s layerwise schedule matters.
Figure 7: Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s insolubility the most.
At the optimal depth (L2), CLaP assigns the highest weights to the aryl bromide and the fused bicyclic ring junction; the 3D branch aggregates the planar $\pi$-surface into a single hydrophobic patch, while the 2D branch captures bromination and ring fusion from connectivity alone. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (a large, contiguous hydrophobic surface) is retained. The outcome matches physical chemistry: expanded aryl surface together with Br increases lipophilicity and contributes essentially no hydrogen-bond donors/acceptors, yielding minimal polar counterbalance and driving $\log S$ more negative (lower aqueous solubility). Under-peeling spreads attribution across the scaffold (shortcuts), and over-peeling smears weights and reintroduces residual context. CLaP effectively captures a general rule in physical chemistry: addition of halogen groups increases the molecule’s overall solubility in organic solvents and decreases the molecule’s overall solubility in aqueous solvents.
C.2 FreeSolv Dataset

### v2

@@ Appendix B Additional Experiments
B.1. Effect of Peeling Hyperparameters
We analyze how the peeling configuration, parameterized by the depth $L$ and the target correlation $\rho_{\max}$, influences performance. Both $L$ and $\rho_{\max}$ are treated as tunable hyperparameters, and we observe that extreme settings can induce under- or over-peeling. Empirically, we find that the optimal configurations of $L$ and $\rho_{\max}$ typically occur at relatively shallow depths, indicating that effective settings can be identified with limited hyperparameter tuning effort.
Peeling depth. We first vary the number of causal blocks $L\in\{3,5,7,9\}$ and observe a U-shaped response in MAE/MSE on CycPeptMPDB in the left panel of Table 3. With too few blocks (L=3L{=}3), the hierarchy under-peels. It lacks capacity to progressively strip context, leaving causal and context signals entangled and hurting accuracy. With too many blocks (L=7,9L{=}7,9), the hierarchy over-peels. The longer optimization path increases variance and can misroute label-relevant patterns into the context stream, degrading generalization. An interior optimum at L=5L{=}5 aligns with a bias–variance view: depth must be large enough to separate signals but not so large that true causal evidence is shaved off together with context.
Table 3. Layers and $\rho_{\max}$ vs. performance on CycPeptMPDB.
Peeling Depth ($L$) Peeling Strength ($\rho_{\max}$)
$L$ MAE MSE $\rho_{\max}$ MAE MSE
3 0.3265 0.1903 0.60 0.3131 0.1741
[…]
7 0.3296 0.2020 0.80 0.3056 0.1644
[…]
Peeling strength. We examine the effect of the target correlation parameter $\rho_{\max}\!\in\!\{0.60,0.70,0.80,0.90\}$ on CycPeptMPDB in the right panel of Table 3. Smaller targets (0.60–0.70) under-peel. The causal branch is not pushed hard enough, so contextual residue persists, and errors rise. A moderate target (0.80) yields the best trade-off. Pursuing near-perfect within-batch correlation (0.90) over-peels. To satisfy an aggressive target, the model exploits batch-coupled features that should remain context, increasing variance and hurting generalization. Appendix A.3 provides additional evidence that an overly large target correlation causes context leakage.
The proposed causal split mechanism generates explicit causal weights for each input component. Each causal block applies node-wise gating on atoms, producing $\alpha_{i}^{(\ell,m)}\!\in\![0,1]$ for atom/group $i$ in modality $m\!\in\!\{\textit{SM},\textit{PE},\textit{GE}\}$ at depth $\ell$. We take the final-layer weight as the causal score, $\pi_{i}^{(m)}=\alpha_{i}^{(L,m)}\in[0,1]$, and render causal saliency maps $\{\pi_{i}^{(m)}\}$ to inspect learned behavior.
Higher $\pi_{i}^{(m)}$ indicates atoms routed through the causal branch at the final layer. Figure 5 compares the under-peeling, optimal-peeling, and over-peeling settings. In the optimal case, saliency concentrates on chemically meaningful moieties, consistent with chemical intuition. At the optimal depth, CLaP concentrates attribution on chemically coherent drivers per task. Across the four datasets, A. ESOL highlights that polar nitro groups and heteroatoms consistently dominate, reflecting their strong role in hydrogen bonding and aqueous solubility (especially in water); B. FreeSolv shows that hydrocarbon segments together with the thio-linker carry the highest weights, aligning with sulfur’s tendency to raise local hydrophobicity and thus overall lipophilicity; and C. Lipo indicates that phenolic OH substituents and ring heteroatoms elevate hydrophilicity while the extended aromatic surface is down-weighted, capturing the trade-off between polar handles and nonpolar bulk consistent with established lipophilicity trends. Finally, in D. CycPeptMPDB dataset, contiguous aromatic and aliphatic side chains form lipophilic patches that promote passive permeation, whereas backbone hydrogen-bond donors and acceptors contribute less, consistent with the desolvation cost of membrane crossing. This pattern reflects the intended peel—the shallow layers under-localize the signal, the deep layers over-peel and smear attribution, and the middle layers isolate chemically relevant substructures that drive the property.
B.2. Additional Ablation Study on Framework Design
Table 4. Additional ablation study on framework design across four datasets. Best indicated in Bold.
Setting ESOL FreeSolv Lipo CycPeptMPDB
MAE MSE MAE MSE MAE MSE MAE MSE
Avg causal layer 0.4798 0.4377 0.9612 1.5843 0.5621 0.5041 0.3227 0.1859
[…]
CLaP 0.4456 0.3583 0.7020 0.8866 0.4672 0.3645 0.3056 0.1644
@@ B.2. Additional Ablation Study on Framework Design
Average causal across layers. This ablation probes the depthwise correlation design. The targets $\{\rho_{\ell}\}$ increase with depth and, together with $\mathcal{L}_{\mathrm{corr}}$ and $\mathcal{L}_{\mathrm{mono}}$ (Eqs. equation 5, equation 6), train the deepest causal block to become progressively more label–relevant. Here we replace the final-layer causal output with a uniform average of $\{{\bm{c}}^{(\ell)}\}_{\ell=1}^{L}$. Given $\mathrm{corr}_{1}\!\leq\!\cdots\!\leq\!\mathrm{corr}_{L}$ under $\mathcal{L}_{\mathrm{corr}}+\mathcal{L}_{\mathrm{mono}}$, averaging $\bar{c}=\tfrac{1}{L}\sum_{\ell}{\bm{c}}^{(\ell)}$ mixes in low-correlation early representations. For centered variables, $\mathrm{Corr}(\bar{c},y)\leq\mathrm{Corr}({\bm{c}}^{(L)},y)$, with strict inequality unless all ${\bm{c}}^{(\ell)}$ are colinear with ${\bm{c}}^{(L)}$. Thus using ${\bm{c}}^{(L)}$ is theoretically favored.
@@
In our experiments, averaging across layers undermines depthwise specialization and mixes early-layer representations that retain substantial context information into the final predictor, contaminating the causal signal and degrading performance (e.g. MSE: $0.164\!\to\!0.185$). These results validate the design: the last causal layer concentrates the causal signal, while earlier layers serve as progressively coarser precursors.
@@
Together, the correlation schedule shapes what each depth should capture; the monotonicity term enforces how the hierarchy evolves; the context branch determines where non-causal residuals go.
@@ C.1. ESOL Dataset
For Figure 7, at the optimal depth (L2), CLaP assigns the highest weights to chloro substituents and adjacent aryl carbons that cluster via short Cl–C/Cl–Cl distances in 3D; the fused readout preserves this through-space signal while the 2D branch highlights halogenated positions from connectivity alone. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (a single hydrophobic patch) is retained. The outcome matches physical chemistry: contiguous chlorination increases non-polar surface and reduces hydrogen-bonding capacity, driving logS lower. Under-peeling spreads attribution across the ring (shortcuts), and over-peeling reintroduces residual context, illustrating why CLaP’s layerwise schedule matters.
Figure 7. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s water affinity the most.
For Figure 8, at the optimal depth (L2), CLaP assigns the highest weights to the aryl bromide and the fused bicyclic ring junction; the 3D branch aggregates the planar $\pi$-surface into a single hydrophobic patch, while the 2D branch captures bromination and ring fusion from connectivity alone. This depth-wise focus is exactly the intended “peel”: batch-coupled context is removed, and label-relevant structure (a large, contiguous hydrophobic surface) is retained. The outcome matches physical chemistry: expanded aryl surface together with Br increases lipophilicity and contributes essentially no hydrogen-bond donors/acceptors, yielding minimal polar counterbalance and driving $\log S$ more negative (lower aqueous solubility). Under-peeling spreads attribution across the scaffold (shortcuts), and over-peeling smears weights and reintroduces residual context. CLaP effectively captures a general rule in physical chemistry: addition of halogen groups increases the molecule’s overall solubility in organic solvents and decreases the molecule’s overall solubility in aqueous solvents.
Figure 8. Layer-wise attribution maps across input modalities. Red colored atoms are attributed with affecting the molecule’s insolubility the most.
C.2. FreeSolv Dataset